OpenStudy (jaden_davis):

A bag contains 3 red marbles, 2 blue marbles, and 5 green marbles. What is the probability of randomly selecting a red marble then, without replacing it, randomly selecting a green marble?

OpenStudy (chrishogan):

A 50% chance.

OpenStudy (chrishogan):

Did it give answer choices?

OpenStudy (.mehek.):

First we set up the fractions : Selecting red: \(\dfrac{3}{10}\) Since we're not replacing, the total goes down by one Selecting green: \(\dfrac{5}{9}\) Now multiply the fractions \(\dfrac{3}{10}*\dfrac{5}{9}\) = ?

OpenStudy (jaden_davis):

thank you

OpenStudy (.mehek.):

You're welcome :)